OpenStudy (anonymous):

Can someone please help me rationalize this expression and state any restrictions?

OpenStudy (anonymous):

\[\frac{ 12-x-x^2 }{ x^2-8x+15}\]

OpenStudy (***[isuru]***):

first factorize the both numerator and denominator...

OpenStudy (***[isuru]***):

and this fraction is already in the rationalized form... u have to rationalize a fraction when there r surds in the denominator ... so all u have to do is mention the restrictions

OpenStudy (***[isuru]***):

\[\frac{ 12 - x - x^{2} }{ x^{2} - 8x +15 } = \frac{ 12 -4x + 3x -x^{2} }{ x^{2} - 3x - 5x +15 }\]

OpenStudy (mathstudent55):

@***[ISURU]*** What are "surds" ?

OpenStudy (***[isuru]***):

\[\frac{ 4( 3-x) + x( 3 - x)}{ x( x-3) -5(x -3) } \]